# A fast neural model for Bayesian Mixed-Effects Regression

### v1

@@ Abstract
Hierarchical data with multiple observations per group is ubiquitous in empirical sciences and is often analyzed using mixed-effects regression. In such models, Bayesian inference gives an estimate of uncertainty but is analytically intractable and requires costly approximation using Markov Chain Monte Carlo (MCMC) methods. Neural posterior estimation shifts the bulk of computation from inference time to pre-training time, amortizing over simulated datasets with known ground truth targets. We propose metabeta, a transformer-based neural network model for Bayesian mixed-effects regression. Using simulated and real data, we show that it reaches stable and comparable performance to MCMC-based parameter estimation at a fraction of the usually required time.
@@ 1 Introduction
Much of the data we work with has a hierarchical structure that naturally clusters into subgroups. When predicting the efficacy of a drug, for example, there may be subpopulation-specific effects in addition to effects on the population as a whole. When building a movie recommendation system, some films may be universally popular, yet individual preferences still matter. When looking at plant growth, the same fertilizer may perform well in one field but poorly in another due to local conditions. These challenges can be addressed using mixed-effects models, which provide a principled framework for capturing both overall trends (fixed effects) and group-specific deviations (random effects). Mixed-effects models have been widely adopted across disciplines – including ecology, psychology, and education – and are by now considered a standard approach for analyzing hierarchical data (Gelman & Hill, 2007; Harrison et al., 2018; Gordon, 2019; Yu et al., 2022).
In many such applications, we would like to estimate the parameters of a mixed-effects model in a Bayesian manner, enabling the incorporation of prior knowledge and the explicit quantification of uncertainty (Figueroa-Zúñiga et al., 2013; Gelman et al., 2013). However, closed-form solutions are generally unavailable even for the simplest cases, necessitating computationally expensive approximate inference methods such as Markov Chain Monte Carlo (MCMC, Metropolis et al., 1953). From a practitioner’s perspective, this is undesirable as MCMC often entails prohibitively long inference times, even for moderately sized datasets.
In this work, we introduce metabeta, a probabilistic transformer-based neural network model, that is designed to efficiently approximate Bayesian inference for mixed-effects regression. It is trained via neural posterior estimation (Rezende & Mohamed, 2015; Gordon et al., 2018; Wildberger et al., 2023; Hollmann et al., 2025), amortizing computation costs over many simulated hierarchical datasets with available ground truth parameters. We demonstrate that metabeta achieves accuracy comparable to Hamiltonian Monte Carlo (HMC), which is the gold-standard MCMC method for Bayesian mixed-effects regression (Neal, 2011; Betancourt, 2018; Bürkner, 2018; Capretto et al., 2022). Importantly, our model reduces inference time by orders of magnitude, thereby greatly broadening the range of feasible applications for Bayesian mixed-effects regression. To further facilitate metabeta’s adoption for rapid deployment and plug-and-play compatibility, we provide open-source Python code for our implementation and plan to release a package with pretrained models that integrates seamlessly with existing analysis pipelines (Bürkner, 2018; Abril-Pla et al., 2023).
@@ 1.1 Related Work
Many methods for neural posterior estimation (NPE) have been proposed in recent years: TabPFN (Müller et al., 2021; Hollmann et al., 2025) is a transformer-based model that efficiently estimates a one-dimensional histogram-like posterior over outcomes $\mathbf{y}$. Its training data is simulated using random graphs, granting it intricate dependence-structure and excellent generalization for non-linear probabilistic prediction. Rezende & Mohamed (2015) pioneered the use of conditional normalizing flows (Papamakarios et al., 2021; Kobyzev et al., 2021) for NPE. Together with Gordon et al. (2018), they laid the groundwork for BayesFlow, a seminal framework combining transformer-based models and normalizing flows (Radev et al., 2020; 2023). In this framework, model parameters are sampled from a fixed prior, observations are sampled from a likelihood, and a neural network learns to map the observations to a joint posterior over model parameters. BayesFlow has been extended to hierarchical Bayesian models with separate networks for each level of hierarchy (Habermann et al., 2024), and to non-linear mixed-effects models optimized for cell biology and pharmacology (Arruda et al., 2023). In both cases, the priors are fixed and a new model has to be trained if a different prior is desired. This off-loads the amortization process to potential end-users, which at best nullifies the runtime advantage of NPE for practical purposes.
Our contribution consists of three aspects: (1) Our model is trained on simulations with varying data ranges and varying parameter priors, explicitly incorporating prior information into posterior estimation; (2) it deploys post-hoc refinements of posterior means and credible intervals using importance sampling (Tokdar & Kass, 2010) and conformal prediction (Vovk et al., 2022); (3) we aim to release a trained version of our model for data practitioners.
@@ 2 Methods
We briefly formalize mixed-effects regression (Section 2.1) and define a synthetic distribution over hierarchical datasets representative of scenarios practitioners care about (Section 2.2). We then present a neural network architecture that takes an entire dataset and priors as inputs and returns posterior distributions over all regression parameters (Section 2.3). This model is trained on synthetic datasets with available ground truth to perform accurate posterior inference (Section 2.4). In a final post-training step, we refine the model’s outputs using importance sampling and conformal prediction (Section 2.5). All our code is implemented in PyTorch 2.7.1 (Paszke et al., 2019) and openly available at https://github.com/adkipnis/metabeta.
@@ 2.1 Mixed-Effects Regression
Mixed-effects regression extends traditional regression by explicitly accounting for within-group dependency in hierarchical data (Gelman & Hill, 2007; Brown, 2021; Fahrmeir et al., 2013). To model this dependency, mixed-effects regression distinguishes between two effect types:
  • Fixed effects $\bm{\beta}\in\mathbb{R}^{d}$ capture the general, group-independent relation between predictor variables $\mathbf{X}_{i}\in\mathbb{R}^{n_{i}\times d}$ and the regression output variable $\mathbf{y}_{i}\in\mathbb{R}^{n_{i}}$.
  • Random effects $\bm{\alpha}_{i}\in\mathbb{R}^{q}$ capture additional, group-specific variations for $q\leq d$ predictors. For each group $i=1,\ldots,m$, we treat $\bm{\alpha}_{i}$ as samples from $\mathcal{N}_{q}(\mathbf{0},\mathbf{S})$.
@@
$\mathbf{y}_{i}=\mathbf{X}_{i}\bm{\beta}+\mathbf{Z}_{i}\bm{\alpha}_{i}+\bm{\varepsilon}_{i}\;,\vskip 14.22636pt$ (1)
@@
with independent additive noise $\bm{\varepsilon}_{i}\sim\mathcal{N}_{n_{i}}(\mathbf{0},\sigma_{\varepsilon}^{2}\mathbf{I}_{n_{i}})$. The random effect predictor matrix $\mathbf{Z}_{i}$ is typically a submatrix of $\mathbf{X}_{i}$. Note, that the $n_{i}$ observations are conditionally independent given some fixed $\bm{\alpha}_{i}$ but marginally dependent over $\bm{\alpha}_{i}$:
$\mathbf{y}_{i}|\bm{\alpha}_{i}\sim\mathcal{N}_{n_{i}}(\mathbf{X}_{i}\bm{\beta}+\mathbf{Z}_{i}\bm{\alpha}_{i},\;\sigma_{\varepsilon}^{2}\mathbf{I}_{n_{i}})\quad\implies\quad\mathbf{y}_{i}\sim\mathcal{N}_{n_{i}}(\mathbf{X}_{i}\bm{\beta},\;\mathbf{Z}_{i}\mathbf{S}\mathbf{Z}_{i}^{\top}+\sigma_{\varepsilon}^{2}\mathbf{I}_{n_{i}}).$
@@
The goal of Bayesian mixed-effects modeling is to obtain posteriors for all unobserved global ($\bm{\beta},\sigma_{\varepsilon}^{2},\mathbf{S}$) and local ($\bm{\alpha}_{i}$) regression parameters, conditioned on the observed predictors, outcomes and priors of the global parameters.
2.2 Data Simulation and Preprocessing
@@
To train our neural posterior estimator, we simulate hierarchically structured datasets using PyTorch as shown in Figure 1A.
Priors: For each dataset, we sample multi-dimensional priors – that is, for $q$ random effect variances $\sigma_{i}$ there are $q$ half-normal priors $\mathcal{HN}(\tau^{2}_{\sigma_{i}})$.
@@
Regression parameters: (1) $q+1$ variance parameters are sampled from half-normal distributions, the first $q$ being the random effect variances, and the last one the noise variance. (2) Then, $m\times q$ random effect vectors are sampled from a diagonal Gaussian, and $d$ fixed effects are sampled from another diagonal Gaussian. (3) Independent noise is sampled from a half-normal distribution.
Observations: The predictors $\mathbf{x}_{ij}$ are sampled independently from several distributions – including normal, Student-$t$, continuous uniform, Bernoulli, negative binomial, and scaled Beta distributions – with randomly chosen parameters. A correlation matrix is sampled from a Lewandowski-Kurowicka-Joe distribution and its lower triangular is multiplied with the predictors to induce correlation structure (Lewandowski et al., 2009). The random effects predictors are set to $\mathbf{z}_{ij}=\mathbf{x}_{ij}$ for $j\leq q$ and 0 otherwise. Predictors and parameters are passed through equation 1 and noise is added to generate outcomes $\mathbf{y_{i}}$ for each group $i$.
Further details on the simulation procedure can be found in Appendix A. We simulate separate mixed-effects datasets for training, validation, and testing. For the latter, we use PyMC (Abril-Pla et al., 2023) to estimate all posteriors with HMC (with the default configuration: $4$ chains, $1000$ tuning iterations, and $1000$ draws). For the HMC model specification, we provide the true priors and the generative model used for simulation. We occasionally encountered divergence and strong outliers for some HMC chains, which may strongly affect performance statistics. For a fair comparison, we choose the chain with the fewest outliers identified by the median absolute deviation statistic (MAD, Hampel, 1974; Leys et al., 2013) for any given dataset.
@@
Figure 1: (A) Dataset Simulation. Given a set of priors, we sample regression parameters and noise in a cascading way. Predictors are sampled from various distributions for training and from real datasets for testing, and outcomes are generated according to equation 1. (B) Model Pipeline. Observed data are summarized locally (per group) and globally (across groups). During training, the posterior networks learn the forward mapping from the true regression parameters to a simple multivariate base distribution, conditioned on the respective summaries and priors. During inference, we draw k samples from the base distribution, and apply the implicitly learned backward mapping to them, approximating sampling from the unknown target posterior. (C) Example Posteriors. Kernel density estimates from the posterior samples of metabeta (MB) and Hamiltonian Monte Carlo (HMC) on a toy dataset. (D) Compute Time. For test sets with $d=5$, $q=1$, $m\leq 30$ and $n_{i}\leq 70$, our model takes several orders of magnitude less time to compute in comparison to HMC. Computation time was measured on a MacBook Air M2 with 24GB of RAM.
@@ 2.3 Model Architecture
The model architecture takes inspiration from BayesFlow (Radev et al., 2020; Habermann et al., 2024) and TabPFN (Hollmann et al., 2025) and has two main parts: (1) a summary network that computes a maximally informative dataset statistic over observations, and (2) a posterior network that uses the summary and priors to propose a joint posterior over regression parameters. Both are trained end-to-end. Since mixed-effect datasets are hierarchically structured, we use two summary and posterior networks, one for the global parameters (fixed effects and variance parameters) and one for the local parameters (group-specific random effects). The training and inference pipeline is visualized in Figure 1B. Data preprocessing is detailed in Appendix B and Appendix C.
@@
Datasets vary in the number of groups and observations per group. A summary network $f_{\Sigma}$ extracts information for the posterior by pooling over all instances in a dataset. Since the data is structured hierarchically, it needs to be summarized accordingly over all exchangeable instances: In a first step, we pool over observations per group, generating $m$ local summaries $\mathbf{s}_{1},\ldots,\mathbf{s}_{m}$. In a second step, we pool the local summaries over groups, generating a global summary $\mathbf{s}$. For summarization, we opted for a set transformer (Lee et al., 2019). Our implementation consists of multiple transformer encoder blocks (Vaswani et al., 2017), followed by averaging over the resulting sequence of transformer outputs. This yields the important property of permutation invariance, i.e. the summary stays the same regardless of the input ordering along the sequence dimension. The local and global summary network both consist of 4 transformer encoder blocks with 128 units, equally large feedforward layers, 8 attention heads, 1% dropout and GELU activations (Hendrycks & Gimpel, 2023).
@@
Posterior networks $f_{\Pi}$ take the dataset summaries and priors as inputs and propose a joint posterior for a set of parameters. Inference on global and local parameters is separated as proposed by Heinrich et al. (2023). Inference on global parameters $\bm{\vartheta}=(\bm{\beta}$, $\mathbf{S}$, $\sigma_{\varepsilon}^{2})$ is conditioned on the global summary and the parameter priors. Inference on local variables ($\bm{\alpha}_{i}$) is conditioned on the separate local summaries and the global parameters (the true ones during training, and the inferred ones during validation). We opted for a normalizing flow (Papamakarios et al., 2021) as our posterior network:
A normalizing flow learns an invertible mapping from a $d$-dimensional random variable $\mathbf{z}_{n}$ with a complex distribution to a $d$-dimensional random variable $\mathbf{z_{0}}$ with a regular distribution (e.g. a multivariate normal). The flow consists of a finite composition $T$ of continuously differentiable and invertible transforms $T_{i}$ with triangular Jacobians, $T=T_{n}\circ\cdots\circ T_{1}$. For some random variable $\mathbf{z}_{0}\sim\mathcal{N}_{d}(\mathbf{0},\mathbf{I})$, we model $T(\mathbf{z}_{n})=\mathbf{z}_{0}\iff T^{-1}(\mathbf{z}_{0})=\mathbf{z}_{n}$ with $p_{n}(\mathbf{z}_{n})=p_{0}(\mathbf{z}_{0})\,\begin{vmatrix}\det J_{T}(\mathbf{z}_{0})\end{vmatrix}$. Each invertible transform $T_{i}$ is parameterized by a neural network that takes part of the current hidden state $\mathbf{z}_{t}$ and the summary $\mathbf{s}$ as inputs. Because of their efficiency, we opted for conditional affine coupling as our normalizing flow architecture (Dinh et al., 2014; 2017). For the base distribution we use a diagonal multivariate location-scale t distribution with learnable parameters for each dimension (Alexanderson & Henter, 2020). For both posterior networks, we use 4 affine coupling flow blocks parameterized by MLPs with three 128-unit feedforward layers, skip connections (He et al., 2016), 1% dropout and ReLU activations.

### v2

@@ Abstract
Hierarchical data with multiple observations per group is ubiquitous in empirical sciences and is often analyzed using mixed-effects regression. In such models, Bayesian inference gives an estimate of uncertainty but is analytically intractable and requires costly approximation using Markov Chain Monte Carlo (MCMC) methods. Neural posterior estimation shifts the bulk of computation from inference time to pre-training time, amortizing over simulated datasets with known ground truth targets. We propose metabeta, a neural network model for Bayesian mixed-effects regression. Using simulated and real data, we show that it reaches stable and comparable performance to MCMC-based parameter estimation at a fraction of the usually required time, enabling new use cases for Bayesian mixed-effects modeling.
Neural Posterior Estimation, Bayesian Mixed-Effects Regression, ICML
@@ 1 Introduction
Much of the data we work with has a hierarchical structure that naturally clusters into subgroups. When predicting the efficacy of a drug, for example, there may be subpopulation-specific effects in addition to effects on the population as a whole. When building a movie recommendation system, some films may be universally popular, yet individual preferences still matter. When looking at plant growth, the same fertilizer may perform well in one field but poorly in another due to local conditions. These challenges can be addressed using mixed-effects models, which provide a principled framework for capturing both overall trends (fixed effects) and group-specific deviations (random effects). Mixed-effects models have been widely adopted across disciplines – including ecology, psychology, and education – and are by now considered the standard approach for statistical analysis of hierarchical data (Gelman and Hill, 2007; Harrison et al., 2018; Gordon, 2019; Yu et al., 2022). In many such applications, we would like to estimate the parameters of a mixed-effects model in a Bayesian manner, enabling the incorporation of prior knowledge and the explicit quantification of uncertainty (Figueroa-Zúñiga et al., 2013; Gelman et al., 2013). However, closed-form solutions are generally unavailable even for the simplest cases, necessitating computationally expensive approximate inference methods such as Markov Chain Monte Carlo (MCMC, Metropolis et al., 1953). From a practitioner’s perspective, this is undesirable, as MCMC typically requires significant runtime and can be brittle with regards to hyper-parameters and design choices (Papaspiliopoulos et al., 2007; Betancourt and Girolami, 2013; Roy, 2019).
In this work, we introduce metabeta, a probabilistic neural network model, that is designed to efficiently approximate Bayesian inference for mixed-effects regression. It is trained via neural posterior estimation (Gordon et al., 2018; Wildberger et al., 2023; Hollmann et al., 2025), amortizing computation costs over many simulated hierarchical datasets with available ground truth parameters. We demonstrate that metabeta achieves accuracy comparable to Hamiltonian Monte Carlo (HMC), which is the gold-standard MCMC method for Bayesian mixed-effects regression (Neal, 2011; Bürkner, 2018; Capretto et al., 2022). Importantly, our model reduces inference time by orders of magnitude, thereby greatly broadening the range of feasible applications for Bayesian mixed-effects regression. To further facilitate metabeta’s adoption for rapid deployment and plug-and-play compatibility, our model is implemented in PyTorch (Paszke et al., 2019) and our code is open-sourced (github.com/adkipnis/metabeta).
@@ 1.1 Related Work
Neural posterior estimation (NPE) — the simulation-based amortization of a neural network posterior — has a long and well-established history. Early work by Papamakarios and Murray (2016) introduced neural conditional density estimators for directly approximating posteriors from simulations. This approach was extended by Lueckmann et al. (2017), who incorporated importance weighting to enable sequential refinement of posterior approximations, and by Greenberg et al. (2019), who proposed automatic posterior transformation, increasing flexibility in proposal adaptation and posterior modeling. These methods form the core foundations of amortized simulation-based Bayesian inference.
Rezende and Mohamed (2015) pioneered the use of conditional normalizing flows (Papamakarios et al., 2021; Kobyzev et al., 2021) for amortized inference. Together with Gordon et al. (2018), they laid the groundwork for BayesFlow (Radev et al., 2020, 2023), which introduced a practical workflow for globally amortized Bayesian inference using summary encoders and normalizing flows. Subsequent extensions adapted BayesFlow to hierarchical Bayesian models (Habermann et al., 2024) and to non-linear mixed-effects models for cell biology and pharmacology (Arruda et al., 2023). In both cases, the prior distribution is fixed at training time, requiring retraining whenever a user wishes to change the prior. This off-loads the amortization process to potential end-users, which strongly diminishes the runtime advantage of NPE for practical purposes.
More recently, transformer-based architectures have emerged as a distinct line of research for amortized Bayesian inference. For instance Distribution Transformers (Whittle et al., 2025) represent prior and posterior as Gaussian Mixture Models whose parameters are mapped by transformers. A thorough comparison of transformer-based NPE methods was recently conducted by Mittal et al. (2025). These works demonstrate that transformer-based NPE can adapt efficiently to varying priors and heterogeneous datasets. However, they have not been tailored specifically to mixed-effects regression, and explicit incorporation of priors in NPE remains an active field of research.
@@ 2 Methods
We briefly formalize mixed-effects regression (Section 2.1) and define a synthetic distribution over hierarchical datasets representative of scenarios practitioners care about (Section 2.2). We then present a neural network architecture that takes an entire dataset and priors as inputs and returns posterior distributions over all regression parameters (Section 2.3). This model is trained on synthetic datasets with available ground truth to perform accurate posterior inference (Section 2.4). In a final post-training step, we refine the model’s outputs using importance sampling and conformal prediction (Section 2.5).
@@ 2.1 Mixed-Effects Regression
Mixed-effects regression extends traditional regression by explicitly accounting for within-group dependency in hierarchical data (Gelman and Hill, 2007; Fahrmeir et al., 2013; Brown, 2021). To model this dependency, mixed-effects regression distinguishes between two effect types:
  • Fixed effects $\boldsymbol{\beta}\in\mathbb{R}^{d}$ capture the general, group-independent relation between predictor variables $\mathbf{X}_{i}\in\mathbb{R}^{n_{i}\times d}$ and the regression output variable $\mathbf{y}_{i}\in\mathbb{R}^{n_{i}}$.
  • Random effects $\boldsymbol{\alpha}_{i}\in\mathbb{R}^{q}$ capture additional, group-specific variations for $q\leq d$ predictors. For each group $i=1,\ldots,m$, we treat $\boldsymbol{\alpha}_{i}$ as samples from $\mathcal{N}_{q}(\mathbf{0},\mathbf{S})$ [^1] For practical purposes $\mathbf{S}\in\mathbb{R}^{q\times q}$ is often constrained to be diagonal, and we adopt this convention.
@@
$\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{Z}_{i}\boldsymbol{\alpha}_{i}+\boldsymbol{\varepsilon}_{i}\;,$ (1)
@@
with independent additive noise $\boldsymbol{\varepsilon}_{i}\sim\mathcal{N}_{n_{i}}(\mathbf{0},\sigma_{\varepsilon}^{2}\mathbf{I}_{n_{i}})$. The random effect predictor matrix $\mathbf{Z}_{i}$ is typically a submatrix of $\mathbf{X}_{i}$. Note, that the $n_{i}$ observations are conditionally independent given some fixed $\boldsymbol{\alpha}_{i}$ but marginally dependent over $\boldsymbol{\alpha}_{i}$:
$\mathbf{y}_{i}|\boldsymbol{\alpha}_{i}\sim\mathcal{N}_{n_{i}}(\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{Z}_{i}\boldsymbol{\alpha}_{i},\;\sigma_{\varepsilon}^{2}\mathbf{I}_{n_{i}})$
$\implies\mathbf{y}_{i}\sim\mathcal{N}_{n_{i}}(\mathbf{X}_{i}\boldsymbol{\beta},\;\mathbf{Z}_{i}\mathbf{S}\mathbf{Z}_{i}^{\top}+\sigma_{\varepsilon}^{2}\mathbf{I}_{n_{i}}).$
@@
The goal of Bayesian mixed-effects modeling is to obtain posteriors for all unobserved global ($\boldsymbol{\beta},\sigma_{\varepsilon}^{2},\mathbf{S}$) and local ($\boldsymbol{\alpha}_{i}$) regression parameters, conditioned on the observed predictors, outcomes and priors of the global parameters.
2.2 Data Simulation
@@
To train our neural posterior estimator, we simulate hierarchically structured datasets as shown in Figure 1A.
Priors: For each dataset, we sample hyper-parameters that specify each multidimensional prior. That is, for $d$ fixed effects, we first sample a $d$-dimensional prior, from which the $d$ fixed effects are sampled later.
@@
Regression parameters: We use the default prior families of Bambi (Capretto et al., 2022), a user-friendly python package for Bayesian modeling with a PyMC backend (Abril-Pla et al., 2023): (1) $q$ random effect variance parameters are sampled from half-normal distributions, (2) then, $m\times q$ random effect vectors are sampled from a diagonal Gaussian using these variance parameters. Independently, $d$ fixed effects are sampled from another diagonal Gaussian. (3) Noise variance is sampled from a half-$t$-distribution, and then independent noise is sampled from a normal distribution with this variance.
@@
Figure 1: (A) Dataset Simulation. Given a set of priors, we sample regression parameters and noise in a cascading way. The random effects covariance matrix is set to $\mathbf{S}=\mathrm{diag}(\sigma_{k})_{k=1,\ldots,q}$. Outcomes are generated according to Equation 1. (B) Model Pipeline. Observations are summarized locally (per group) and globally (across groups). During training, the posterior networks learn the conditional forward mapping (true regression parameters $\to$ multivariate base distribution). During inference, we sample from the base distribution, and apply the backward mapping, approximating sampling from the unknown target posterior.
Observations: The predictors $\mathbf{x}_{ij}$ are sampled from two sources: Synthetic distributions and real datasets from the PMLB (Romano et al., 2021) and SRM (Lichtenberg and Şimşek, 2017) benchmarks. The random effects predictors are set to $\mathbf{z}_{ij}=\mathbf{x}_{ij}$ for $j\leq q$ and 0 otherwise. Predictors are standardized and passed through Equation 1 with the regression parameters and noise to generate outcomes $\mathbf{y_{i}}$ for each group $i$. Further details on the simulation procedure can be found in Appendix A.
Baselines: For the test sets, we use Bambi on top of PyMC to estimate all posteriors using Hamiltonian Monte Carlo (HMC) and Variational Inference (VI, Blei et al., 2017), which is the standard computationally cheaper alternative to MCMC. For HMC we use the No-U-Turn sampler (Hoffman and Gelman, 2011) with four chains ($2500$ tuning iterations and $1000$ posterior draws per chain). For VI we use automatic differentiation variational inference (Kucukelbir et al., 2016) with the ADAM optimizer (Kingma and Ba, 2017) ($5\times 10^{-3}$ learning rate, $5\times 10^{5}$ training iterations and $4000$ draws). We supply the true priors and the generative model used for simulation, so the model specification follows the ground truth. For a fair comparison, we exclude datasets with divergent MCMC samples from the test set.
@@ 2.3 Model Architecture
The model architecture takes inspiration from BayesFlow (Radev et al., 2020; Habermann et al., 2024) and TabPFN (Hollmann et al., 2025) and has two main components: (1) a summary network that computes a maximally informative dataset statistic over observations, and (2) a posterior network that uses the summary and priors to propose a joint posterior over regression parameters. Both are trained end-to-end. Since mixed-effect datasets are hierarchically structured, we use two summary and posterior networks, one for the global parameters (fixed effects and variance parameters) and one for the local parameters (group-specific random effects). The training and inference pipeline is visualized in Figure 1B. Data preprocessing is detailed in Appendix B and Appendix C.
@@
Datasets vary in the number of groups and observations per group. A summary network $f_{\Sigma}$ extracts information for the posterior by pooling over all instances in a dataset. Since the data is structured hierarchically, it needs to be summarized accordingly over all exchangeable instances: In a first step, we pool over observations per group, generating $m$ local summaries $\mathbf{s}_{1},\ldots,\mathbf{s}_{m}$. In a second step, we pool the local summaries over groups, generating a global summary $\mathbf{s}$. For summarization, we opted for a set transformer (Lee et al., 2019). Our implementation consists of multiple transformer encoder blocks (Vaswani et al., 2017), followed by averaging over the resulting sequence of transformer outputs. This yields the important property of permutation invariance, i.e. the summary stays the same regardless of the input ordering along the sequence dimension. The local and global summary network both consist of 3 transformer encoder blocks with 128 units, equally large feedforward layers, 8 attention heads, 1% dropout and GELU activations (Hendrycks and Gimpel, 2023).
@@
Posterior networks $f_{\Pi}$ take the dataset summaries and priors as inputs and propose a joint posterior for a set of parameters. Inference on global and local parameters is separated for hierarchical NPE (Rodrigues et al., 2021; Heinrich et al., 2023; Habermann et al., 2024). Inference on global parameters $\boldsymbol{\vartheta}=(\boldsymbol{\beta}$, $\mathbf{S}$, $\sigma_{\varepsilon}^{2})$ is conditioned on the global summary and the parameter priors. Inference on local variables ($\boldsymbol{\alpha}_{i}$) is conditioned on the separate local summaries and the global parameters (the true ones during training, and the inferred ones during validation). We opted for a normalizing flow (Papamakarios et al., 2021) as our posterior network:
A normalizing flow learns an invertible mapping from a $d$-dimensional random variable $\mathbf{z}_{n}$ with a challenging distribution to a $d$-dimensional random variable $\mathbf{z_{0}}$ with a well-behaved distribution (e.g. a multivariate normal). The flow consists of a finite composition $T$ of continuously differentiable and invertible transforms $T_{i}$ with triangular Jacobians, $T=T_{n}\circ\cdots\circ T_{1}$. For some random variable $\mathbf{z}_{0}\sim\mathcal{N}_{d}(\mathbf{0},\mathbf{I})$, we model $T(\mathbf{z}_{n})=\mathbf{z}_{0}\iff T^{-1}(\mathbf{z}_{0})=\mathbf{z}_{n}$ with $p_{n}(\mathbf{z}_{n})=p_{0}(\mathbf{z}_{0})\,\begin{vmatrix}\det J_{T}(\mathbf{z}_{0})\end{vmatrix}$. Each invertible transform $T_{i}$ is parameterized by a neural network that takes part of the current hidden state $\mathbf{z}_{t}$ and the summary $\mathbf{s}$ as inputs. Because of their efficiency, we opted for conditional affine coupling as our normalizing flow architecture (Dinh et al., 2014, 2017). For the base distribution we use a diagonal multivariate location-scale t distribution with learnable parameters for each dimension (Alexanderson and Henter, 2020). For both posterior networks, we use 8 affine coupling flow blocks parameterized by MLPs with three 256-unit feedforward layers, skip connections (He et al., 2016), 1% dropout and ReLU activations.